Mary buys 100 candies for $7. If yips cost 5 cents each, Gups 6 cents each and pips 7 cents each, how many more pips than yips did Mary buy?
A 50 B 60 C 70 D 100 E none of the above

Answer by richwmiller(17219)

6g+7p+5y=700
g+p+y=100
p=y+100
D)100